# Harmonic covers of skeleta

### v1

@@ 2. Background on logarithmic regularity
In this section we recall some well-known properties of log-regular log-schemes [Kat94]. Some properties of monoids, fans and polyhedral geometry are recalled in Appendix A. Experienced readers may wish to skip to Section 3 and refer back as needed. Comprehensive references on log-geometry include [Ogu18, Kat89, Kat94, Niz06, GR18, Tem24].
@@
Any log-regular schemes is normal [Kat94, §4.1] and a fs log-scheme log-smooth111 Recall that a morphism $f^{\dagger}:X^{\dagger}\longrightarrow Y^{\dagger}$ of coherent log-schemes is called log-smooth if it is locally of finite type and formally log-smooth (defined via infinitesimal liftings of log-thickenings in [Ogu18, §IV.3.1]). Log-smooth morphisms are preserved by composition and fs base change [Ogu18, IV.3.2]. In practice one uses Kato’s criterion [Kat89, §3.5]or [Ogu18, §3.1.13]: if $\phi:P\to Q$ is a coherent chart at $x\in X$ such that #$\#\ker\phi^{\mathrm{gp}}$ and #$\#\left(\mathrm{coker}\phi^{\mathrm{gp}}\right)_{\mathrm{tor}}$ are invertible in $k(x)$, and $X\to Y\times_{\mathrm{Spec}\,\mathbb{Z}[P]}\mathrm{Spec}\,\mathbb{Z}[Q]$ is smooth (in the usual sense), then $f$ is log-smooth at $x$; and analogously for log-étale morphisms. over a log-regular scheme is log-regular [Kat94, §8.2].
@@
For example suppose that $X$ is a regular integral scheme of dimension $d$ and $D_{X}$ defines a (reduced) snc divisor on $X$ [Sta, 0CBN]; recall this means that at every point $x\in X$ there exists a system of regular parameters $t_{1},\dots,t_{d}\in\mathscr{O}_{X,x}$ such that $D_{X,x}$ is defined by $\prod_{i=1}^{r}t_{i}=0$ in $\mathscr{O}_{X,x}$ for some $r\leq d$. Then $X^{\dagger}=X(\mathrm{log}D_{X})$ is log-regular: indeed, near $x$ we have a sharp fs chart $\mathbb{N}^{r}\to\mathscr{O}_{X}:(n_{1},\dots,n_{r})\mapsto t_{1}^{n_{1}}\cdots t_{r}^{n_{r}}$ so $\mathscr{C}_{X,x}\cong\mathbb{N}^{r}$ and $\mathrm{rank}(x)=r$, and $\mathcal{I}_{X,x}$ is generated by $t_{1},\dots,t_{r}$ with $\mathrm{codim}V(\mathcal{I}_{X,x})=r$.
In fact we have a converse: if $X^{\dagger}$ is a log-regular log-scheme then by 2.5 we have $X^{\dagger}=X(\mathrm{log}D_{X}))$ and if the underlying scheme $X$ is regular then $D_{X}$ is a snc divisor: indeed, at each $x\in X$, the quotient of the regular ring $\mathscr{O}_{X,x}$ by $\mathcal{I}_{X,x}$ is regular, which implies that $\mathcal{I}_{X,x}$ is generated by regular parameters by [Sta, 00NR].
@@
called the characteristic morphism of $X^{\dagger}$, which sends a point $x$ to the prime ideal defined by $I_{X^{\dagger},x}$ in $\mathrm{Spec}\,\mathscr{O}_{X,x}$, it is open and in fact initial among all open maps of $(X,\mathscr{M}_{X}^{\sharp})$ to a sharp monoidal space (see op.cit. or [Kat94, 10.2]), this gives a categorical222informally, one can think of $\pi$ as a categorical quotient, though unfortunately this does not make sense since fans and log-schemes live in different categories. It is plausible one can remedy this by using Artin stacks instead of Kato fans – see [ACM+15] and the work of Ulirsch [Uli18]. definition of $F(X^{\dagger})$. For every morphism $f:X^{\dagger}\to Y^{\dagger}$ of log-regular log-schemes we have an induced map $F(f):F(X^{\dagger})\to F(Y^{\dagger})$ of fans, obtained via the composition $F(X^{\dagger})\to(X,\mathscr{C}_{X})\to(Y,\mathscr{C}_{Y})\to F(Y^{\dagger})$.
@@
Let $X^{\dagger}$ be log-regular log-scheme. Each point of $F(X^{\dagger})$ is locally closed, therefore since $\pi=\pi_{X^{\dagger}}:(X,\mathscr{M}_{X})\to F(X^{\dagger})$ is open, the fibers of $\pi$ define a stratification of $X$ into irreducible locally closed subsets, here called the log-stratification, which by [BM19, Lemma 2.2.3] coincides with the stratification defined by the support $D_{X^{\dagger}}$. Hence the points of $F(X^{\dagger})$ parametrise log-strata. For a point $x\in F(X^{\dagger})$ we denote by $\pi^{-1}{x}=E^{\circ}_{x}$ the associated log-stratum and $E_{x}=\overline{\{x\}}$ for its schematic closure. We call $E_{x}^{\circ}$ (resp. $E_{x}$) the log-stratum (resp. closed log-stratum) associated to $x$. Note that the log-stratification refines the rank stratification of 2.2.
@@
Omitted, this is the heart result of [Kat94, §9-10], also see [GR18, §6.6] for a more detailed proof.
For later reference we recall how to construct $(X^{\prime})^{\dagger}$: suppose that the proper subdivision $F^{\prime}\to F$ is affine-locally given by $\mathrm{Spec}\,\phi:\mathrm{Spec}\,P^{\prime}\to\mathrm{Spec}\,P$ for some morphism $\phi:P\to P^{\prime}$ of sharp fs monoids. Then $X^{\prime}$ is locally constructed as
@@
The proof of ‘̀log-crepant”, see LABEL:log-crepant. in the in the sense that This will play a role in
@@
Recall that locally fs (Kato) fans correspond to cone complexes, see Section A.26. Given a log-regular log-scheme $X^{\dagger}$ we write $\Sigma(X^{\dagger})=\Sigma(F(X^{\dagger})$ for the cone complex associated to $F(X^{\dagger})$ – its construction is recalled in A.25. In summary, it is the geometrical realisation of $F({X}^{\dagger})$ defined by glueing, for all $x\in F({X}^{\dagger})$, the dual cones
@@
along the cospecialisation maps as face embeddings, see (A.24). For shortness, we will also write $C(x)\coloneqq\sigma_{\mathscr{C}_{F({X}^{\dagger}),x}}\subset\Sigma({X}^{\dagger})$ for the cone associated to $x\in F({X}^{\dagger})$. Note that
@@
where $F_{0}=F(\Sigma_{0})$ is the Kato fan associated to $\Sigma_{0}$. Let $\Sigma^{\prime}_{0}$ denote the cone complex associated to $F_{0}^{\prime}$. and let $\varphi:\Sigma(X^{\prime\dagger})\to\Sigma(X^{\dagger})$ denote the associated morphism of cone complexes. For each cone $\sigma^{\prime}$ of $\Sigma(X^{\dagger})$, the cone $\varphi(\sigma^{\prime})$ is a finite union $\cup\sigma_{i}$ of cones $\sigma_{i}$ in $\Sigma_{0}$, because $\Sigma_{0}\to\Sigma(X^{\dagger})$ is a proper subdivision. By construction $\Sigma^{\prime}_{0}$ is obtained by glueing the cones $\phi^{-1}\sigma_{i}$. So by definition $\Sigma_{0}^{\prime}\to\Sigma(X^{\prime\dagger})$ is a proper subdivision (Definition A.10). Therefore so is $F_{0}^{\prime}\to F^{\prime}$.
@@ 3. Balanced covers of dual complexes
Recall the following running assumptions, see Conventions 1.32. Throughout $k^{\circ}$ denotes a discretely valued field with a choice of uniformiser $\varpi$. The scheme $S=\mathrm{Spec}\,k^{\circ}$ is endowed with the standard log-structure $S^{\dagger}$ given by the closed point.
@@
We define a toroidal $S$-scheme as a log-scheme ${X}^{\dagger}=({X},\mathscr{M}_{{X}})$ such that ${X}$ is an integral flat separated finite type $S$-scheme.
@@
denotes a finite morphism of toroidal $S$-schemes, which will also call a toroidal cover. Note this implies $f^{-1}D_{{X}}\subset D_{X^{\prime}}$.
@@
Let us show that in nature, one can often encounter morphisms $f:X^{\prime\dagger}\to{X}^{\dagger}$ as in 3.4; we give four situations below. Let ${X}^{\dagger}$ be a toroidal $S$-scheme.
@@
    Suppose ${X}^{\dagger}\to S^{\dagger}$ is log-smooth, and $S^{\prime\dagger}\to S^{\dagger}$ is a finite morphism of log-schemes, then $X^{\prime\dagger}\coloneqq{X}^{\dagger}\times^{\mathrm{fs}}S^{\prime\dagger}\to{X}^{\dagger},$ is log-smooth, and hence $X^{\prime\dagger}$ is log-regular. In particular the underlying scheme $X^{\prime}$ is normal, and so coincides with the normalisation of the base change ${X}_{S^{\prime}}={X}\times_{S}S^{\prime}$.
@@
    Suppose ${X}$ is regular of dimension $2$ and write $U={X}\setminus D_{{X}}$. Suppose $U^{\prime}\to U$ is a finite étale cover which is cyclic of degree $p$. Then by Abhyankar’s theorem on $p$-cyclic covers of Jungian domains – see Kato’s comments in [Kat94, §12.4] and the references cited there – it follows that there exists a modification $g:{X}_{0}\to{X}$ such that the following conditions (i)-(iii) hold:
@@
    Note that if (i)-(iii) hold then $f$ is a finite morphism of $2$-dimensional toroidal $S$-schemes (the morphism $f$ is finite because ${X}_{0}$ is excellent).
@@
We mention that Kato [Kat94, §12.4.1] has asked the question whether the assertion of 3.5 (4) remains true for any finite étale cover $U^{\prime}\to U$ and in arbitrary dimension. By the existence of toroidal resolutions (2.13) and Lemma 2.16, we may additionally require ${X}_{0}$ to be regular (as Kato does) without changing the outcome of this question. To the author’s knowledge Kato’s question is still wide open.
@@
Throughout the remainder of the article $f:{X}^{\dagger}\to X^{\prime\dagger}$ is as in Assumption 3.4. The constructions of Section 2.7 yield a morphism of Kato fans $F(X^{\prime\dagger})\to F({X}^{\dagger})$. Since the morphism ${X}^{\dagger}\to S^{\dagger}$ provides structural maps $\mathbb{N}\to\mathscr{C}_{F({X}^{\dagger}),x}$, sending $1$ to the class of the uniformiser $[\varpi]\in\mathscr{C}_{F({X}^{\dagger}),x}$, we can and will view $F({X}^{\dagger})$ as a fan over $\mathrm{Spec}\,\mathbb{N}$, and $F(X^{\prime\dagger})\to F({X}^{\dagger})$ as a morphism of fans over $\mathrm{Spec}\,\mathbb{N}$. To make the notation somewhat less heavy, we will also write
@@
Note that by definition, $M_{x}$ is a lattice and $\mathrm{rank}(M_{x})=\mathrm{rank}(x)$.
@@
The map $\mathscr{C}_{F({X}^{\dagger}),x}\to\mathscr{C}_{F(X^{\prime\dagger}),x^{\prime}}$ is local morphism of sharp monoids, which by sharpness is the same as injective monoid morphism (A.18). By the assumption that $x^{\prime}\in F(X^{\prime\dagger})$ and $f:X^{\prime}\to{X}$ is finite, we have a finite morphism of log-strata
@@
The dual complex $\mathcal{D}({X}^{\dagger})$ of ${X}^{\dagger}$ is defined as the $\mathbb{Z}$-PL subspace of $\Sigma({X}^{\dagger})$ defined by glueing the polyhedrons
@@
by restricting the face maps of $\Sigma({X}^{\dagger})$. This defines a functor from the category of log-regular log-schemes over $S^{\dagger}$ to $\mathbb{Z}$-PL complexes. Throughout we denote by
@@
the induced morphism of $\mathbb{Z}$-PL complexes. Note that the following are equivalent:
@@
Moreover, $\sigma_{x}$ is bounded if and only if the closed stratum $E_{x}$ associated to $x$ (see 2.8) is an intersection of vertical components (components contained in ${X}_{s}$) of $D=D_{{X}}$.
@@
Informally speaking, the points of $\sigma_{x}$ are “$k$-valuations” on the monoid $\mathscr{C}_{F({X}^{\dagger}),x}$, in the sense that they define monoid morphisms $\mathscr{C}_{F({X}^{\dagger},x}\to\mathbb{R}_{\geq 0}$ which are compatible with the normalised valuation of $k$.
@@
Let $f:X^{\prime\dagger}\to{X}^{\dagger}$ be a toroidal cover. Let $y^{\prime}\in F(X^{\prime\dagger})$, write $y=f(y)$ and let $x\in\overline{\{y\}}$ be a specialisation of $y$. Then
@@
By Lemma 2.9 we have a diagram with exact rows
@@
The desired formula now follows from Equations (3.13) and (3.14) and Lemma 3.15 below. ∎
@@
Note that $\sigma$ is strictly convex if and only if $\sigma^{\prime}$ is so. In that case $(M_{\sigma}^{+})^{\mathrm{gp}}=M_{\sigma}$ where $M_{\sigma}^{+}$ denotes the monoid of non-negative $\mathbb{Z}$-affine functions on $\sigma$ (see A.32).
@@
Let $\sigma,\tau$ be polyhedrons of a PL-space $\Sigma$. Then we write $\sigma>\tau$ if and only if $\tau$ is a proper face of $\sigma$. We will also call $\sigma$ an adjacent polyhedron of $\tau$. Suppose $n=\dim\Sigma\geq 1$. A facet is defined as a $n$-dimensional polyhedron and a ridge is defined as a $(n-1)$-dimensional polyhedron. By a facet-ridge pair $(\sigma,\tau)$ we mean a pair of a facet $\sigma\in\Sigma(n)$ and a ridge $\tau\in\Sigma(n-1)$ such that $\sigma>\tau$.
@@
The balancing notion of 3.20 behaves well with proper subdivisions in the following sense. If $\Sigma^{\prime}\to\Sigma$ is a balanced finite cover of $\mathbb{Z}$-PL complexes and $\Sigma_{0}\to\Sigma$ is a proper subdivision, then $\Sigma^{\prime}_{0}=\Sigma^{\prime}\times_{\Sigma}\Sigma_{0}\to\Sigma_{0}$ is a balanced finite cover of $\mathbb{Z}$-PL complexes and $\Sigma^{\prime}_{0}\to\Sigma^{\prime}$ is a proper subdivision.
Theorem 3.24.
@@
It remains to prove that $\deg\phi=\deg f$. For any vertex $v\in\mathcal{D}({X}^{\dagger})(0)$ (a 0-dimensional polyhedron), Lemma 3.15 and equation (3.13) (applied to the generic point of $X^{\prime}$) yield the formula
@@
Another application of Proposition 3.11 then shows that the last expression equals $\deg\phi$. ∎
@@
A variant of Theorem 3.24 is true without the assumption that $\tilde{k}$ is algebraically closed, but next to the $\mathbb{Z}$-PL structure on $\mathcal{D}({X}^{\dagger})$ one would have to additionally remember the degree $\deg_{\widetilde{k}}{\sigma}\coloneqq\left[\kappa(x):\widetilde{k}\right]$ on each facet $\sigma=\sigma_{x}$ of $\mathcal{D}({X}^{\dagger})$, and change the balancing condition to $\deg_{\tau^{\prime}}\phi=\sum\deg_{\sigma^{\prime}}\phi$ where $\deg_{\sigma^{\prime}}\phi\coloneqq[M_{\sigma^{\prime}}:M_{\sigma}]\deg_{\widetilde{k}}{\sigma^{\prime}}/\deg_{\widetilde{k}}{\sigma}$ (in accordance with Proposition 3.11).
4. Metrisation of $\mathbb{Z}$-PL spaces

### v2

@@ 2. Background on logarithmic regularity
In this section we give a short background exposition of log-regular log-schemes [Kat94]. Some properties of monoids, fans and polyhedral geometry are recalled in Appendix A. Experienced readers may wish to skip to Section 3 and refer back as needed. Comprehensive references on log-geometry include [Ogu18, Kat89, Kat94, Niz06, GR18, Tem24].
@@
Any log-regular scheme is normal [Kat94, §4.1] and a fs log-scheme log-smooth222 Recall that a morphism $f^{\dagger}:X^{\dagger}\longrightarrow Y^{\dagger}$ of coherent log-schemes is called log-smooth if it is locally of finite type and formally log-smooth (defined via infinitesimal liftings of log-thickenings in [Ogu18, §IV.3.1]). Log-smooth morphisms are preserved by composition and fs base change [Ogu18, IV.3.2]. In practice one uses Kato’s criterion [Kat89, §3.5]or [Ogu18, §3.1.13]: if $\phi:P\to Q$ is a coherent chart at $x\in X$ such that #$\#\ker\phi^{\mathrm{gp}}$ and #$\#\left(\mathrm{coker}\phi^{\mathrm{gp}}\right)_{\mathrm{tor}}$ are invertible in $k(x)$, and $X\to Y\times_{\mathrm{Spec}\,\mathbb{Z}[P]}\mathrm{Spec}\,\mathbb{Z}[Q]$ is smooth (in the usual sense), then $f$ is log-smooth at $x$; and analogously for log-étale morphisms. over a log-regular scheme is log-regular [Kat94, §8.2].
@@
Suppose that $X$ is a regular integral scheme of dimension $d$ and $D_{X}$ defines a (reduced) snc divisor on $X$ [Sta, 0CBN]; recall this means that at every point $x\in X$ there exists a system of regular parameters $t_{1},\dots,t_{d}\in\mathscr{O}_{X,x}$ such that $D_{X,x}$ is defined by $\prod_{i=1}^{r}t_{i}=0$ in $\mathscr{O}_{X,x}$ for some $r\leq d$. Then $X^{\dagger}=X(\mathrm{log}D_{X})$ is log-regular: indeed, near $x$ we have a sharp fs chart $\mathbb{N}^{r}\to\mathscr{O}_{X}:(n_{1},\dots,n_{r})\mapsto t_{1}^{n_{1}}\cdots t_{r}^{n_{r}}$ so $\mathscr{C}_{X,x}\cong\mathbb{N}^{r}$ and $\mathrm{rank}(x)=r$, and $\mathcal{I}_{X,x}$ is generated by $t_{1},\dots,t_{r}$ with $\mathrm{codim}V(\mathcal{I}_{X,x})=r$.
Conversely if $X^{\dagger}$ is a log-regular log-scheme then by 2.5 we have $X^{\dagger}=X(\mathrm{log}D_{X}))$ and if the underlying scheme $X$ is regular then $D_{X}$ is a snc divisor: indeed, at each $x\in X$, the quotient of the regular ring $\mathscr{O}_{X,x}$ by $\mathcal{I}_{X,x}$ is regular, which implies that $\mathcal{I}_{X,x}$ is generated by regular parameters by [Sta, 00NR].
@@
called the characteristic morphism of $X^{\dagger}$, which sends a point $x$ to the prime ideal defined by $I_{X^{\dagger},x}$ in $\mathrm{Spec}\,\mathscr{O}_{X,x}$, it is open and in fact initial among all open maps of $(X,\mathscr{M}_{X}^{\sharp})$ to a sharp monoidal space (see op.cit. or [Kat94, 10.2]), this gives a categorical333informally, one can think of $\pi$ as a categorical quotient, though unfortunately this does not make sense since fans and log-schemes live in different categories. It is plausible one can remedy this by using Artin stacks instead of Kato fans – see the work of Ulirsch [Uli18]. definition of $F(X^{\dagger})$. For every morphism $f:X^{\dagger}\to Y^{\dagger}$ of log-regular log-schemes we have an induced map $F(f):F(X^{\dagger})\to F(Y^{\dagger})$ of fans, obtained via the composition $F(X^{\dagger})\to(X,\mathscr{C}_{X})\to(Y,\mathscr{C}_{Y})\to F(Y^{\dagger})$.
@@
Let $X^{\dagger}$ be log-regular log-scheme. Each point of $F(X^{\dagger})$ is locally closed. Since $\pi=\pi_{X^{\dagger}}:(X,\mathscr{M}_{X})\to F(X^{\dagger})$ is open, the fibers of $\pi$ define a stratification of $X$ into irreducible locally closed subsets, here called the log-stratification, which by [BM19, Lemma 2.2.3] coincides with the stratification defined by the support $D_{X^{\dagger}}$. Hence the points of $F(X^{\dagger})$ parametrise log-strata. For a point $x\in F(X^{\dagger})$ we denote by $\pi^{-1}{x}=E^{\circ}_{x}$ the associated log-stratum and $E_{x}=\overline{\{x\}}$ for its schematic closure. We call $E_{x}^{\circ}$ and $E_{x}$ the log-stratum associated to $x$ and closed log-stratum associated to $x$ respectively. The log-stratification refines the rank stratification of 2.2.
@@
Omitted, this is the main result of [Kat94, §9-10], also see [GR18, §6.6] for a more detailed proof. For later reference we recall how to construct $(X^{\prime})^{\dagger}$: suppose that the proper subdivision $F^{\prime}\to F$ is affine-locally given by $\mathrm{Spec}\,\phi:\mathrm{Spec}\,P^{\prime}\to\mathrm{Spec}\,P$ for some morphism $\phi:P\to P^{\prime}$ of sharp fs monoids. Then $X^{\prime}$ is locally constructed as
@@
Recall that locally fs (Kato) fans correspond to cone complexes, see Section A.26. Given a log-regular log-scheme $X^{\dagger}$ we write $\Sigma(X^{\dagger})=\Sigma(F(X^{\dagger}))$ for the cone complex associated to $F(X^{\dagger})$ – its construction is recalled in A.25. In summary, it is the geometric realisation of $F({X}^{\dagger})$ defined by glueing, for all $x\in F({X}^{\dagger})$, the dual cones
@@
along the cospecialisation maps as face embeddings, see (A.24). For brevity we will also write $C(x)\coloneqq\sigma_{\mathscr{C}_{F({X}^{\dagger}),x}}\subset\Sigma({X}^{\dagger})$ for the cone associated to $x\in F({X}^{\dagger})$. Note that
@@
where $F_{0}=F(\Sigma_{0})$ is the Kato fan associated to $\Sigma_{0}$. Let $\Sigma^{\prime}_{0}$ denote the cone complex associated to $F_{0}^{\prime}$ and let $\varphi:\Sigma(X^{\prime\dagger})\to\Sigma(X^{\dagger})$ denote the associated morphism of cone complexes. For each cone $\sigma^{\prime}$ of $\Sigma(X^{\dagger})$, the cone $\varphi(\sigma^{\prime})$ is a finite union $\cup\sigma_{i}$ of cones $\sigma_{i}$ in $\Sigma_{0}$, because $\Sigma_{0}\to\Sigma(X^{\dagger})$ is a proper subdivision. By construction $\Sigma^{\prime}_{0}$ is obtained by glueing the cones $\phi^{-1}\sigma_{i}$. So by definition (see A.10) $\Sigma_{0}^{\prime}\to\Sigma(X^{\prime\dagger})$ is a proper subdivision and so is $F_{0}^{\prime}\to F^{\prime}$.
@@ 3. Balanced covers of dual complexes
Recall the running assumptions of Conventions 1.31. The scheme $S=\mathrm{Spec}\,k^{\circ}$ is endowed with the standard log-structure $S^{\dagger}$ given by the closed point.
@@
We define a toroidal $S$-scheme as a log-scheme ${X}^{\dagger}=({X},\mathscr{M}_{{X}})$ such that ${X}$ is an integral flat separated finite type $S$-scheme and ${X}^{\dagger}$ is log-regular.
@@
denotes a finite morphism of toroidal $S$-schemes, which we will also call a toroidal cover. Note this implies $f^{-1}D_{{X}}\subset D_{X^{\prime}}$.
@@
To justify our choice of setting in 3.4, let us show that in nature, one can often encounter morphisms $f:X^{\prime\dagger}\to{X}^{\dagger}$ as in 3.4; we give four situations below. Let ${X}^{\dagger}$ be a toroidal $S$-scheme.
@@
    Suppose ${X}^{\dagger}\to S^{\dagger}$ is log-smooth, and $S^{\prime\dagger}\to S^{\dagger}$ is a finite morphism of log-schemes, then $X^{\prime\dagger}\coloneqq{X}^{\dagger}\times^{\mathrm{fs}}S^{\prime\dagger}\to{X}^{\dagger},$ is log-smooth, and hence $X^{\prime\dagger}$ is log-regular. In particular the underlying scheme $X^{\prime}$ is normal and coincides with the normalisation of the base change ${X}_{S^{\prime}}={X}\times_{S}S^{\prime}$.
@@
    Suppose ${X}$ is regular of dimension $2$ and write $U={X}\setminus D_{{X}}$. Suppose $U^{\prime}\to U$ is a finite étale cover which is cyclic of degree $p$. Then by Abhyankar’s theorem [Abh66] on $p$-cyclic covers of Jungian domains – see Kato’s comments in [Kat94, §12.4] and the references cited there – it follows that there exists a modification $g:{X}_{0}\to{X}$ such that the following conditions (i)-(iii) hold:
@@
    If (i)-(iii) hold then $f$ is a finite morphism of $2$-dimensional toroidal $S$-schemes (the morphism $f$ is finite because ${X}_{0}$ is excellent).
@@
We mention that Kato [Kat94, §12.4.1] has asked the question whether the assertion of 3.5 (4) remains true for any finite étale cover $U^{\prime}\to U$ and in arbitrary dimension. By the existence of toroidal resolutions (2.13) and Lemma 2.16, we may additionally require ${X}_{0}$ to be regular (as Kato does) without changing the outcome of this question. As far as the author is aware Kato’s question is still wide open.
@@
For the remainder of the paper $f:{X}^{\dagger}\to X^{\prime\dagger}$ is as in Assumption 3.4, sometimes additional assumptions are imposed and then this is indicated in the beginning of a new section. The constructions of Section 2.7 yield a morphism of Kato fans $F(X^{\prime\dagger})\to F({X}^{\dagger})$. Since the morphism ${X}^{\dagger}\to S^{\dagger}$ provides structural maps $\mathbb{N}\to\mathscr{C}_{F({X}^{\dagger}),x}$, sending $1$ to the class of the uniformiser $[\varpi]\in\mathscr{C}_{F({X}^{\dagger}),x}$, we can and will view $F({X}^{\dagger})$ as a fan over $\mathrm{Spec}\,\mathbb{N}$, and $F(X^{\prime\dagger})\to F({X}^{\dagger})$ as a morphism of fans over $\mathrm{Spec}\,\mathbb{N}$. To make the notation somewhat less heavy, we will also write
@@
By definition, $M_{x}$ is a lattice and $\mathrm{rank}(M_{x})=\mathrm{rank}(x)$.
@@
The map $\mathscr{C}_{F({X}^{\dagger}),x}\to\mathscr{C}_{F(X^{\prime\dagger}),x^{\prime}}$ is local morphism of sharp monoids, which by sharpness is the same as an injective monoid morphism (A.18). By the assumption that $x^{\prime}\in F(X^{\prime\dagger})$ and $f:X^{\prime}\to{X}$ is finite, we have a finite morphism of log-strata
@@
The dual complex $\mathcal{D}({X}^{\dagger})$ of ${X}^{\dagger}$ is defined as the $\mathbb{Z}$-PL subspace of $\Sigma({X}^{\dagger})$ defined by glueing the polyhedra
@@
by restricting the face maps of $\Sigma({X}^{\dagger})$. This defines functor from the category of log-regular log-schemes over $S^{\dagger}$ to $\mathbb{Z}$-PL complexes and thus functor factors through the category of locally fs fans. Throughout we denote by
@@
the induced morphism of $\mathbb{Z}$-PL complexes. The following are equivalent:
@@
Moreover, $\sigma_{x}$ is bounded if and only if the closed stratum $E_{x}$ associated to $x$ (see 2.8) is an intersection of vertical components of $D_{{X}}$, i.e. components of ${X}_{s}$
@@
Informally speaking, the points of $\sigma_{x}$ are “$k$-valuations” on the monoid $\mathscr{C}_{F({X}^{\dagger}),x}$, in the sense that they define monoid morphisms $\mathscr{C}_{F({X}^{\dagger}),x}\to\mathbb{R}_{\geq 0}$ which are compatible with the normalised valuation of $k$.
@@
Let $f:X^{\prime\dagger}\to{X}^{\dagger}$ be a toroidal cover (3.4. Let $y^{\prime}\in F(X^{\prime\dagger})$, write $y=f(y)$ and let $x\in\overline{\{y\}}$ be a specialisation of $y$. Then
@@
By Lemma 2.9 (iv)-(v) we have a diagram with exact rows
@@
The desired formula now follows from Lemma 2.9 (ii), Equations (3.13) and (3.14) and Lemma 3.15 below. ∎
@@
Note that $\sigma$ is strictly convex if and only if $\sigma^{\prime}$ is so. In that case $(\mathscr{C}_{\sigma})^{\mathrm{gp}}=M_{\sigma}$ where $\mathscr{C}_{\sigma}$ denotes the monoid of non-negative $\mathbb{Z}$-affine functions on $\sigma$ (see A.32).
@@
Let $\sigma,\tau$ be polyhedra of a PL-space $\Sigma$. Then we write $\sigma>\tau$ if and only if $\tau$ is a proper face of $\sigma$. We will also call $\sigma$ an adjacent polyhedron of $\tau$. Suppose $n=\dim\Sigma\geq 1$. A facet is defined as a $n$-dimensional polyhedron and a ridge is defined as a $(n-1)$-dimensional polyhedron. By a facet-ridge pair $(\sigma,\tau)$ we mean a pair of a facet $\sigma\in\Sigma(n)$ and a ridge $\tau\in\Sigma(n-1)$ such that $\sigma>\tau$.
@@
The balancing condition of 3.20 behaves well with respect to proper subdivisions in the following sense. If $\Sigma^{\prime}\to\Sigma$ is a balanced finite cover of $\mathbb{Z}$-PL complexes and $\Sigma_{0}\to\Sigma$ is a proper subdivision, then $\Sigma^{\prime}_{0}=\Sigma^{\prime}\times_{\Sigma}\Sigma_{0}\to\Sigma_{0}$ is a balanced finite cover of $\mathbb{Z}$-PL complexes and $\Sigma^{\prime}_{0}\to\Sigma^{\prime}$ is a proper subdivision.
Theorem 3.24 (Balancing condition).
@@
It remains to prove that $\deg\phi=\deg f$. For any vertex $v\in\mathcal{D}({X}^{\dagger})(0)$, Lemma 3.15 and equation (3.13) (applied to the generic point of $X^{\prime}$) yield the formula
@@
Proposition 3.11 then shows that the right-hand side equals $\deg\phi$. ∎
@@
The assumption that $\tilde{k}$ is algebraically closed is not essential in the proof of Theorem 3.24. If we omit this hypothesis then one additionally decorates each facet $\sigma=\sigma_{x}$ of $\mathcal{D}({X}^{\dagger})$ with the degree $\deg_{\widetilde{k}}{\sigma}\coloneqq\left[\kappa(x):\widetilde{k}\right]$ and the balancing condition is changed to $\deg_{\tau^{\prime}}\phi=\sum\deg_{\sigma^{\prime}}\phi$ where $\deg_{\sigma^{\prime}}\phi\coloneqq[M_{\sigma^{\prime}}:M_{\sigma}]\deg_{\widetilde{k}}{\sigma^{\prime}}/\deg_{\widetilde{k}}{\sigma}$, in accordance with Proposition 3.11.
4. Metrisation of $\mathbb{Z}$-PL complexes